OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y = one halfx + 3.

OpenStudy (anonymous):

slopes of perpendicular line m*m1=-1 where m is slope of given line slope of the given line m=1/2 slope of perpendicular line =-2 point slope form y-)-1)=-2(x-(-3)) y+1=-2x-6 2x+y+7=0

OpenStudy (anonymous):

y = -2x - 7 y = -2x +5 y = 2x + 5 y = 2x - 7

OpenStudy (anonymous):

A